# Bayesianize Fuzziness in the Statistical Analysis of Fuzzy Data

### v1

@@ 1 Introduction
The increasing complexity of data in modern contexts is fundamentally transforming the landscape of analysis and modeling, driving the need for more sophisticated approaches and innovative methodologies. This shift, often described as the transition from the curse to the blessing of dimensionality [12], is driven by the increasing prevalence of unstructured and semi-structured data sources. However, this complexity encompasses more than just the variety and volume of data. Instead, it arises from challenges such as unconventional data structures – like networks and trees – and the need to address diverse sources of uncertainty, including sampling variability, measurement errors, systematic biases, and instrumental distortions [22]. Even within more traditional frameworks, complexities persist. A notable example is the use of rating scales to measure subjective attitudes, opinions, or beliefs regarding specific objects. Though longstanding and practical, these scales are influenced by cognitive, affective, and contextual factors inherent to the response process [49]. Consider, for instance, the evaluation of sensory attributes, such as the overall experience of tasting a sparkling wine using a 7-point hedonic scale ranging from “completely dislike” to “completely like”. In this scenario, raters rely on long-term memory, activate affective and cognitive components, and integrate them to form and express their final opinions. This decision-making process can introduce uncertainty at various stages, as conflicting cognitive and emotional demands influence the final response. Consequently, the recorded response may not fully capture the underlying experience. To address such complexities, fuzzy rating scales have been developed and employed successfully in several applicative contexts [10, 4, 43]. In this setting, fuzzy mathematics provides a valuable framework to represent this type of uncertainty as epistemic imprecision or fuzziness alongside the traditional randomness that drives the sampling process [9]. This approach allows for the simultaneous accommodation of randomness and fuzziness, enhancing the precision and reliability of statistical models. By clarifying ambiguities in empirical data, we can enhance the representation of uncertainty, leading to more robust and interpretable insights amidst increasing data complexity.
@@
Despite the availability of numerous fuzzy statistical tools designed to handle fuzzy data, a consistent and comprehensive framework for their statistical analysis is still missing. In particular, a general methodology comparable to the Generalized Linear Models (GLMs) framework in traditional statistics has yet to be developed. Although these methods are firmly grounded in measure-theoretic principles [32, 19], their practical use is limited by the absence of general operational parametric probability models for fuzzy random variables. Furthermore, the lack of a unified inferential framework that provides broadly applicable asymptotic results adds to the challenges of their implementation [21].
@@
Based on the general premise that fuzzy data arise from a process analogous to statistical coarsening [20, 7, 37], this contribution introduces a comprehensive method that incorporates the epistemic mechanism assumed to underlie the generation of fuzzy data. The proposed approach utilizes a conditional probabilistic framework to connect the observed characteristics of fuzzy numbers (e.g., mode or spread) to the true, yet unobserved, statistical model underlying the random sample prior to fuzzification. Estimation and inference are performed using the Gibbs sampler-based approach, wherein the full conditional distribution is approximated through sampling from a quadratic approximation of the target posterior distribution. Notably, our solution adopts a regression-based framework similar to GLMs, allowing for the incorporation of covariates into the modeling process. As this work represents the first attempt to provide such a generalization, we restrict our attention to the case of bounded fuzzy variables only, which are frequently encountered in social science studies [46].
@@ 2 Preliminaries
Additionally, a BFN can be represented using its mode $m$ and precision $s$, under the assumption that $lb=0$ and $ub=1$, simplifying the notation. In this case, the membership function becomes
@@
This formulation provides a concise and flexible way to model BFNs, simplifying their application in further analyses. For instance, BFNs are particularly useful for approximating various fuzzy numbers, as they can generate flexible shapes suitable for specific tasks, such as Bayesian analysis.
@@ 3.1 Statement of the problem
Let $\mathbb{Y}=\{Y_{1},\ldots,Y_{n}\}$ denote a set of $n$ independent continuous random variables and let $\mathbf{\widetilde{y}}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})$ be a sample of fuzzy observations. Due to epistemic or post-sampling uncertainty-based processes [6], the vector $\mathbf{\widetilde{y}}$ serves as a blurred representation of the true, yet unobserved, vector of crisp random outcomes $\mathbf{y}$. The goal is to study the joint density $f_{Y_{1},\ldots,Y_{n}}(\mathbf{y};\bm{\theta}_{y})$ in order to make inference about the parameter vector $\bm{\theta}_{y}$, given the fuzzy sample $\mathbf{\widetilde{y}}$. From the epistemic perspective, fuzzy observations are regarded as stochastic outcomes influenced by a combination of non-random and systematic uncertainty, which obscure the true but unknown realizations $\mathbf{y}$. In this contribution, each Beta-type fuzzy observation is characterized by a mode and precision, expressed as $\widetilde{y}_{i}=\{m_{i},s_{i}\}$.
@@ 3.2 Proposed solution
A key feature of the integration process described in (11) is its capacity to avoid the trivial case where the mode $m_{i}$ of the observed fuzzy number equals the latent realization $y_{i}$. While this scenario can occur (e.g., see Fig. 2, second panel, counterclockwise), the proposed modeling approach accommodates the broader case where the true unobserved realization lies within a plausible interval. This interval represents partial information about uncertain data captured through a possibility distribution. Such a scenario reflects a typical case where the observer possesses only approximate knowledge of the data and can specify merely a graded interval of plausible values [11]. Importantly, relying on defuzzified data (e.g., the mode $m_{i}$ or center-of-gravity index) instead of the observed fuzzy sets as input for inference on $\bm{\theta}_{y}$ would lead to biased estimates, particularly for variance parameters, as shown in prior studies [5].
To better understand the rationale behind the proposed conditional sampling schema, particularly the interaction of its three terms in shaping the expectation and variance of the modes of the observed fuzzy numbers, consider the following expressions (assuming $lb=0$ and $ub=1$ for simplicity)
@@
Consider $S\sim\mathcal{G}(s;1,\beta_{s})$ in the shape-scale parameterization for simplicity. It is evident that as $\beta_{s}\to 0$ – a scenario representing high fuzziness in the observed fuzzy numbers – the scaling factor $c\to 1$, leading to $\mathbb{V}(M)<\mathbb{E}[Y]$. Conversely, as $\beta_{s}\to\infty$ – representing progressively reduced fuzziness – the scaling factor $c\to 0$, and $\mathbb{V}(M)\approx\mathbb{V}(Y)$
When $c\geq 0.5$, which occurs at $\beta_{s}=1.639$, the dispersion of the observed modes becomes significantly smaller than that of the unobserved crisp realizations. This happens because $c$ attenuates the variance of $Y$, leading to a phenomenon akin to underdispersion [44]. This underdispersion can introduce bias when estimating $\mathbb{V}(Y)$ based on fuzzy observations. To mitigate this issue, cases where $c\ll 0.5$ are preferable – for instance, at $\beta_{s}=10.121$, $c=0.25$, ensure the observed modes’ variance closely approximates that of the unobserved realizations.
Overall, the random variable $S$, which governs the degree of fuzziness, introduces variability that blurs the relationship between $M$ and $Y$ by attenuating $\mathbb{V}(Y)$. This effect becomes more pronounced with high variability in $S$, diminishing the ability of $M$ to fully represent the variability of $Y$. For completeness, derivations of (12)-(13) are provided in the Appendix.
@@ 3.3.4 The model $f_{Y}(\mathbf{y};\bm{\theta}_{y})$
The conditional model is typically designed for fuzzy bounded data. Within a regression framework, several statistical models can be applied, including the Beta [15], Kumaraswamy [36], Logitnormal [40], and Logbilai [38] distributions. However, the true underlying model, such as the normal distribution, may be unbounded. In such cases, the fuzzification process effectively bounds the true unobserved realizations $\{y_{1},\ldots,y_{n}\}$.
@@ 6 Applications
In this section, we demonstrate the application of the proposed model using a collection of pre-existing datasets. The section is structured into four subsections: the first three assess the model’s external validity by evaluating its ability to reproduce observed fuzzy data through a Posterior Predictive Check (PPC) approach, while the final subsection presents its application in a complete case study.
@@
    Estimate the model parameters $\bm{\theta}_{y}$ using the observed fuzzy dataset $\mathbf{\widetilde{y}}{n\times 1}$ using the approximated Gibbs sampler (see Section 4).
@@
For all datasets, the approximate Gibbs sampler described in Section 4 is applied, running five independent chains with 4,000 samples each, following a burn-in period of 2,000 samples. To maintain simplicity, weakly informative priors are used – ensuring vague yet reasonable constraints on parameter values. Chain convergence is assessed using standard diagnostics, including the Gelman-Rubin statistic ($\widehat{R}$) and Effective Sample Size (ESS) [17]. In cases requiring model comparison, the Widely Applicable Information Criterion (WAIC) is employed [51].
@@ 6.1 Reforestation data (Colubi, 2009)
A fuzzy-valued scale was adopted since a standard categorical scale (e.g., very low, low, medium, high, and very high) failed to capture the nuances of these expert perceptions. The scale’s support was expressed as a percentage, with 0 indicating no quality and 1 representing perfect quality. For our analysis, we used a subsample of $n=200$ fuzzy observations. The three tree species (Betula, Quercus, and Sorbus) were encoded as levels of a categorical variable used to predict $\mathbb{E}[Y]$. The original trapezoidal fuzzy numbers were transformed into Beta-type fuzzy numbers using a procedure designed to minimize information loss from the original fuzzy sets, as described in [5].
@@ 6.2 Cognitive response times (Kim & Bishu, 1998)
The second dataset comes from a study on the cognitive response times of nuclear power plant control room crews during abnormal events [31]. Various performance factors, including stress levels and the potential for misdiagnosis, influence these response times. As noted by Kim and Bishu [31], cognitive times can be effectively represented as fuzzy numbers, capturing the subjective nature of human responses under stress. For this study, the entire sample of $n=75$ units was utilized, along with three continuous predictors to model $\mathbb{E}[Y]$: Inside control room experience, Outside control room experience, and Education. The original triangular fuzzy numbers were transformed into Beta-type ones using the procedure detailed in Section [5].
Considering the nature of the response variable, we followed the Kim & Bishu’s choice to adopt the Lognormal distribution for modeling $f_{Y}(y;\bm{\theta}_{y})=\mathcal{LN}(y;\mu_{i},\sigma^{2})$. In this case, as $Y\in\mathbb{R}^{+}$, bounds for the modes $m_{1},\ldots,m_{n}$ were computed by using 0-cuts of the observed fuzzy numbers (see Section 3.3). All the chains runned here achieved the convergence, with the bulk-ESS ranging from 7772 to 8000 and the tail-ESS from 9113 and 9503. The PPC steps (i)–(iii) were applied $B=500$ times for the chosen model. Figure 5 presents the results for the three statistics under consideration. The posterior predictive analysis suggests an adequate goodness-of-fit for the estimated model. The metrics achieved an adequate level of accuracy (centroids: $\text{CP}=0.72$, $\text{bP}=0.01$; supports: $\text{CP}=0.72$, $\text{bP}=0.01$; fuzziness: $\text{CP}=0.88$, $\text{bP}=0.08$).
@@
Since $Y$ represents a bounded rating variable, the Kumaraswamy distribution $f_{Y}(y;\bm{\theta}_{y})=\mathcal{K}(y;\alpha_{i},\nu)$ was chosen to model the non-fuzzy component, with the following parametrization $\nu=(1+\exp(-\phi))^{-1}$ and $\alpha_{i}=-\log(1-0.5^{\nu})/\log(1+\exp(-\mathbf{X}\bm{\beta}))$. Convergence was achieved for all chains, with bulk-ESS ranging from 7011 to 8101 and tail-ESS from 8897 to 9302. The PPC steps (i)–(iii) were applied $B=500$ times for the chosen model. Figure 7 presents the results for the three evaluated statistics. Overall, the posterior predictive analysis suggests a good fit for the estimated model, with the predictive metrics achieving a satisfactory level of accuracy (centroids: $\text{CP}=0.70$, $\text{bP}=0.01$; supports: $\text{CP}=0.90$, $\text{bP}=0.03$; fuzziness: $\text{CP}=0.89$, $\text{bP}=0.02$).

### v2

@@ 1 Introduction
The increasing complexity of data in modern contexts is fundamentally transforming the landscape of analysis and modeling, driving the need for more sophisticated approaches and innovative methodologies. This shift, often described as the transition from the curse to the blessing of dimensionality [13], is driven by the increasing prevalence of unstructured and semi-structured data sources. However, this complexity encompasses more than just the variety and volume of data. Instead, it arises from challenges such as unconventional data structures – like networks and trees – and the need to address diverse sources of uncertainty, including sampling variability, measurement errors, systematic biases, and instrumental distortions [24]. Even within more traditional frameworks, complexities persist. A notable example is the use of rating scales to measure subjective attitudes, opinions, or beliefs regarding specific objects. Though longstanding and practical, these scales are influenced by cognitive, affective, and contextual factors inherent to the response process [52]. Consider, for instance, the evaluation of sensory attributes, such as the overall experience of tasting a sparkling wine using a 7-point hedonic scale ranging from “completely dislike” to “completely like”. In this scenario, raters rely on long-term memory, activate affective and cognitive components, and integrate them to form and express their final opinions. This decision-making process can introduce uncertainty at various stages, as conflicting cognitive and emotional demands influence the final response. Consequently, the recorded response may not fully capture the underlying experience. To address such complexities, fuzzy rating scales have been developed and employed successfully in several applicative contexts [11, 4, 46]. In this setting, fuzzy mathematics provides a valuable framework to represent this type of uncertainty as epistemic imprecision or fuzziness alongside the traditional randomness that drives the sampling process [10]. This approach allows for the simultaneous accommodation of randomness and fuzziness, enhancing the precision and reliability of statistical models.
However, the use of rating scales is just one example where fuzzy data analysis can be conveniently applied. While it is widely assumed that data from random experiments can be numerically represented, several research contexts yield imprecise data that cannot be adequately expressed on a standard numerical scale. Although some researchers simplify their models to avoid this issue, fuzzy random variables (or random fuzzy numbers) offer a natural way to model both randomness and fuzziness in real-world evaluations. Nevertheless, as in any area of mathematical and statistical modeling, a solid theoretical foundation is crucial. In particular, inference based on random fuzzy numbers needs to be framed within a probabilistic structure, ensuring that methodologies for imprecise data are statistically well-justified.
@@
That said, despite the availability of numerous statistical tools designed to handle fuzzy data, a consistent and comprehensive framework for their statistical analysis is still lacking. In particular, a general methodology comparable to the Generalized Linear Models (GLMs) framework in traditional statistics has yet to be developed. Although the fuzzy methods developed so far are firmly grounded in measure-theoretic principles [34, 20], their practical use is limited by the absence of general operational parametric probability models for fuzzy random variables. Furthermore, the lack of a unified inferential framework that provides broadly applicable asymptotic results adds to the challenges of their implementation [23].
@@
Based on the general premise that fuzzy data arise from a process analogous to statistical coarsening [21, 7, 40], this contribution introduces a comprehensive method that incorporates the epistemic mechanism assumed to underlie the generation of fuzzy data. The proposed approach utilizes a conditional probabilistic framework to connect the observed characteristics of fuzzy numbers (e.g., mode or spread) to the true, yet unobserved, statistical model underlying the random sample prior to fuzzification. Estimation and inference are performed using the Gibbs sampler-based approach, wherein the full conditional distribution is approximated through sampling from a quadratic approximation of the target posterior distribution. Notably, our solution adopts a regression-based framework similar to GLMs, allowing for incorporating covariates into the modeling process. As this work represents the first attempt to provide such a generalization, we restrict our attention to the case of bounded fuzzy variables only, which are frequently encountered in social science studies [49].
The main findings presented in this paper can be summarized as follows:
  1. (i)
    An epistemic-oriented two-stage generation scheme for fuzzy samples is introduced. This approach, closely related to a GLM-like hierarchical Bayesian framework, also enables a model-oriented defuzzification procedure. As a result, the underlying parameters of the statistical model can be estimated even when the data are fuzzy or imprecise and obscure their actual values. Although we focus on the case of bounded fuzzy sets represented by Beta-type fuzzy numbers, the proposed approach can also be generalized to more complex cases involving unbounded cases.
  2. (ii)
    Analytical formulas are provided for the Bayesian formulation of the proposed sampling mechanism. To numerically solve these, we employ a quadratic approximation and a multivariate Skew-Normal distribution-based approximation, both used in conjunction with a Gibbs sampling procedure. Our numerical results show that the quadratic approximation yields low errors and provides high-quality estimates.
  3. (iii)
    To demonstrate the practical usefulness of the overall approach (i.e., the sampling scheme and the accompanying numerical procedures), four case studies based on real-life fuzzy data are analyzed. The performance of the proposed method is assessed using posterior predictive checks based on numerical simulations for each dataset. The final case study additionally involves parameter estimation of the underlying model and a detailed analysis of the associated variables. Finally, the Supplementary Materials provide extended numerical and graphical results for the fuzzy data analysis, including a comprehensive comparison with a selection of state-of-the-art fuzzy regression methods and a detailed and illustrated procedure for assessing model misspecification through posterior predictive analyses.
@@ 2 Preliminaries
Additionally, a BFN can be represented using its mode $m$ and precision $s$. In this case, assuming for simplicity $lb=0$ and $ub=1$, the membership function becomes
@@
This formulation provides a concise and flexible way to model BFNs, simplifying their application in further analyses.
@@ 3.1 Statement of the problem
We start with the general statement of the problem. Let $\mathbb{Y}=\{Y_{1},\ldots,Y_{n}\}$ denote a set of $n$ independent continuous random variables and let $\mathbf{\widetilde{y}}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})$ be a sample of fuzzy observations. Due to epistemic or post-sampling uncertainty-based processes [6], the vector $\mathbf{\widetilde{y}}$ serves as a blurred representation of the true, yet unobserved, vector of crisp random outcomes $\mathbf{y}$. The goal is to study the joint density $f_{Y_{1},\ldots,Y_{n}}(\mathbf{y};\bm{\theta}_{y})$ in order to make inference about the parameter vector $\bm{\theta}_{y}$, given the fuzzy sample $\mathbf{\widetilde{y}}$. From the epistemic perspective, fuzzy observations are regarded as stochastic outcomes influenced by a combination of non-random and systematic uncertainty, which obscure the true but unknown realizations $\mathbf{y}$.
In this contribution, we apply a specific type of blur to the true values $\mathbf{y}$, motivated by the epistemic interpretation of fuzzy random variables discussed above. We assume that the fuzzy observations $\mathbf{\widetilde{y}}$ are described by Beta-type fuzzy numbers (see Sect. 3.2), so that each Beta-type fuzzy observation can be conveniently characterized by a mode and a precision parameter, expressed as $\widetilde{y}_{i}=\{m_{i},s_{i}\}$.
@@ 3.2 Proposed solution
A key feature of the integration process described in (11) is its capacity to avoid the trivial case where the mode $m_{i}$ of the observed fuzzy number equals the latent realization $y_{i}$. While this scenario can occur (e.g., see Fig. 2, second panel, counterclockwise), the proposed modeling approach accommodates the broader case where the true unobserved realization lies within a plausible interval. This interval represents partial information about uncertain data captured through a possibility distribution. Such a scenario reflects a typical case where the observer possesses only approximate knowledge of the data and can specify merely a graded interval of plausible values [12]. Note that, relying on defuzzified data (e.g., the mode $m_{i}$ or center-of-gravity index) instead of the observed fuzzy sets as input for inference on $\bm{\theta}_{y}$ would lead to biased estimates, particularly for variance parameters, as shown in prior studies [5].
To better understand the rationale behind the proposed conditional sampling schema, particularly the interaction of its three terms in shaping the expectation and variance of the modes of the observed fuzzy numbers, consider as follows (assuming $lb=0$ and $ub=1$ for simplicity):
@@
Consider $S\sim\mathcal{G}(s;1,\beta_{s})$ in the shape-scale parameterization for simplicity. It is evident that as $\beta_{s}\to 0$ – a scenario representing high fuzziness in the observed fuzzy numbers – the scaling factor $c\to 1$, leading to $\mathbb{V}\textrm{ar}\left[M\right]<\mathbb{E}\left[Y\right]$. Conversely, as $\beta_{s}\to\infty$ – representing progressively reduced fuzziness – the scaling factor $c\to 0$, and $\mathbb{V}\textrm{ar}\left[M\right]\approx\mathbb{V}\textrm{ar}\left[Y\right]$
When $c\geq 0.5$, which occurs at $\beta_{s}=1.639$, the dispersion of the observed modes becomes significantly smaller than that of the unobserved crisp realizations. This happens because $c$ attenuates the variance of $Y$, leading to a phenomenon akin to underdispersion [47]. This underdispersion can introduce bias when estimating $\mathbb{V}\textrm{ar}\left[Y\right]$ based on fuzzy observations. To mitigate this issue, cases where $c\ll 0.5$ are preferable – for instance, at $\beta_{s}=10.121$, $c=0.25$, ensure the observed modes’ variance closely approximates its counterpart for the unobserved realizations.
Overall, the random variable $S$, which governs the degree of fuzziness, introduces variability that blurs the relationship between $M$ and $Y$ by attenuating $\mathbb{V}\textrm{ar}\left[Y\right]$. This effect becomes more pronounced with high variability in $S$, diminishing the ability of $M$ to fully represent the variability of $Y$. For completeness, derivations of (12)-(13) are provided in the Appendix.
@@ 3.3.4 The model $f_{Y}(\mathbf{y};\bm{\theta}_{y})$
The conditional model is typically designed for fuzzy bounded data. In this case, several statistical models can be applied, including the Beta [16], Kumaraswamy [39], Logitnormal [43], and Logbilai [41] distributions. However, when the true underlying model is unbounded, such as for the Gaussian case, the fuzzification process still bounds the actual unobserved realizations $\{y_{1},\ldots,y_{n}\}$ within the set of most plausible values.
@@
3.3.5 Assessing model misspecifications
One can rely on posterior predictive checks and the examination of marginal posterior distributions to assess whether the assumed two-stage fuzzification mechanism adequately captures the data-generating process. In particular, model misspecifications – such as those arising from more complex blurring mechanisms involving additional stages – can be detected by evaluating discrepancies between observed and model-implied predictions. Key indicators include poor coverage of the supports and fuzziness, increased posterior variance, and systematic bias in predictive summary statistics. These diagnostics offer insight into whether the model assumptions are violated, especially in the presence of structurally biased fuzzy data, such as $\delta$-shifted observations that allow for fuzzy data not containing the true unobserved realizations. Further details and examples are provided in Section 3 of the Supplementary Materials.
3.3.6 Model implications in real contexts
It should be highlighted that our proposal is suitable for data collected in contexts where fuzziness reflects partial knowledge, vague measurement, or perceptual ambiguity – such as expert assessments, rating tasks, and decision support systems. However, this proposal does not aim to model the internal cognitive mechanisms behind fuzzification. Instead, more simply, it offers a statistical approximation of epistemic processes, allowing practitioners to analyze imprecise data without assuming that respondents consciously implement the underlying probabilistic model. Because of this, existing fuzzy rating procedures, like those proposed by [11, 4], naturally align with the assumptions of the conditional model and may be seen as compatible backend mechanisms underlying the observed fuzzy data.
That said, in this context, the core of a fuzzy number $\tilde{A}_{1}^{(i)}$ is identified with the mode $m_{i}$, representing the value most compatible with the underlying latent variable $Y_{i}$, such as a rater’s true but unobserved judgment. The support $\tilde{A}_{0}^{(i)}$ is governed by the spread parameter $s_{i}$, which controls the decay of the membership function around the mode. Therefore, the core captures the most plausible value, while the support reflects a broader range of less precise, yet still compatible, values.
@@ 6 Applications
In this section, we illustrate the application of the proposed model using a collection of pre-existing datasets. The section is structured into four subsections: the first three assess the model’s external validity through a Posterior Predictive Check (PPC) approach, while the final subsection presents a complete case study.
@@
    Estimate the model parameters $\bm{\theta}_{y}$ using the observed fuzzy dataset $\mathbf{\widetilde{y}}_{n\times 1}$ using the approximated Gibbs sampler (see Section 4).
@@
For all datasets, the approximate Gibbs sampler described in Section 4 is applied, running five independent chains with 4000 samples each, following a burn-in period of 2000 samples. Weakly informative priors are used, ensuring vague yet reasonable constraints on parameter values. Chain convergence is assessed using standard diagnostics, including the Gelman-Rubin statistic ($\widehat{R}$) and Effective Sample Size (ESS) [18]. In cases requiring model comparison, the Widely Applicable Information Criterion (WAIC) is employed [54].
@@ 6.1 Reforestation data (Colubi, 2009)
A fuzzy-valued scale was adopted since a standard categorical scale failed to capture the nuances of these expert perceptions. The scale’s support was expressed as a percentage, with 0 indicating no quality and 1 representing perfect quality. For our analysis, we used a subsample of $n=200$ fuzzy observations. The three tree species (Betula, Quercus, and Sorbus) were encoded as levels of a categorical variable used to predict $\mathbb{E}[Y]$. The original trapezoidal fuzzy numbers were transformed into Beta-type fuzzy numbers using a procedure designed to minimize information loss from the original fuzzy sets, as described in [5].
@@ 6.2 Cognitive response times (Kim & Bishu, 1998)
The second dataset comes from a study on the cognitive response times of nuclear power plant control room crews during abnormal events [33]. Various performance factors, including stress levels and the potential for misdiagnosis, influence these response times. As noted by Kim and Bishu [33], cognitive times can be effectively represented as fuzzy numbers, capturing the subjective nature of human responses under stress. For this study, the entire sample of $n=75$ units was utilized, along with three continuous predictors to model $\mathbb{E}[Y]$: Inside control room experience ($X_{1}$), Outside control room experience ($X_{1}$), and Education ($X_{3}$). The original triangular fuzzy numbers were transformed into Beta-type ones using the procedure detailed in [5].
Considering the nature of the response variable, we followed the Kim & Bishu’s choice to adopt the Lognormal distribution for modeling $f_{Y}(y;\bm{\theta}_{y})=\mathcal{LN}(y;\mu_{i},\sigma^{2})$. In this case, as $Y\in\mathbb{R}^{+}$, bounds for the modes $m_{1},\ldots,m_{n}$ were computed by using 0-cuts of the observed fuzzy numbers (see Section 3.3). All the chains simulated here converged, with the bulk-ESS ranging from 7772 to 8000 and the tail-ESS from 9113 and 9503. The PPC steps (i)–(iii) were applied $B=500$ times for the chosen model. Figure 5 presents the results for the three statistics under consideration. The posterior predictive analysis suggests a marginally adequate goodness-of-fit for the estimated model (centroids: $\text{CP}=0.99$, $\text{bP}=0.03$; supports: $\text{CP}=0.80$, $\text{bP}=0.18$; fuzziness: $\text{CP}=0.90$, $\text{bP}=0.10$).
@@
Since $Y$ represents a bounded rating variable, the Kumaraswamy distribution $f_{Y}(y;\bm{\theta}_{y})=\mathcal{K}(y;\alpha_{i},\nu)$ was chosen to model the non-fuzzy component, with the following parametrization $\nu=(1+\exp(-\phi))^{-1}$ and $\alpha_{i}=-\log(1-0.5^{\nu})/\log(1+\exp(-\mathbf{X}\bm{\beta}))$. Convergence was achieved for all chains, with bulk-ESS ranging from 7011 to 8101 and tail-ESS from 8897 to 9302. The PPC steps (i)–(iii) were applied $B=500$ times for the chosen model. Figure 7 presents the results for the three evaluated statistics. Overall, the posterior predictive analysis suggests a good fit for the estimated model, with the predictive metrics achieving a satisfactory level of accuracy (centroids: $\text{CP}=0.99$, $\text{bP}=0.04$; supports: $\text{CP}=0.90$, $\text{bP}=0.03$; fuzziness: $\text{CP}=0.90$, $\text{bP}=0.02$).